OpenStudy (anonymous):

tanx-1/over/tanx+1=1-cotx/over/1+cotx

OpenStudy (anonymous):

lets call \(\cos(x)=a\) and \(\sin(x)=b\) then the left hand side is \[\frac{\frac{b}{a}-1}{\frac{b}{a}+1}\] multiply top and bottom by \(a\) to simplify the complex fraction and get \[\frac{b-a}{b+a}\] the right hand side will be \[\frac{1-\frac{a}{b}}{1+\frac{a}{b}}\] multiply top and bottom by \(b\) and get \(\frac{b-a}{b+a}\) same as the left this really has nothing to do with trig, it is just algebra

OpenStudy (anonymous):

multiply the top by cot(x)

OpenStudy (anonymous):

and multiply the bottom by cot(x)

OpenStudy (anonymous):

can anyone kindly tell me does this editor use Latex scripting